OpenStudy (anonymous):

1/3x+5/6=13/14 /this is another clearing fraction eqation

jhonyy9 (jhonyy9):

1/3x +5/6 =13/14 2+x ---- =13/14 6x 7(2+x)=39x --- 14+7x=39x --- 14=32x --- x=14/32 =7/16